OpenStudy (anonymous):

Use both the first and second derivative tests to show that 3x^2-6x+1 has a relative minimum at x=1.

OpenStudy (freckles):

have you found the derivative of the given expression?

OpenStudy (freckles):

first derivative =0 will give us the possible mins and maxs

OpenStudy (freckles):

or at least where they occur of course

OpenStudy (anonymous):

i derived it and got 6x-6 for the first and 6 for the second derivavtive is thatright?

OpenStudy (freckles):

that sounds great.

OpenStudy (freckles):

6x-6=0 gives us?

OpenStudy (anonymous):

yeah so does that mean 1 is a critical point?

OpenStudy (freckles):

yes

OpenStudy (freckles):

since since f''>0 then that implies with have what kinda extrema at x=1?

OpenStudy (freckles):

Remember I told you earlier what f''>0 implies

OpenStudy (anonymous):

relative minimum?

OpenStudy (freckles):

yes we have a rel min at x=1

OpenStudy (freckles):

\[f(x)=3x^2-6x+1 \\ f'(x)=6x-6 \\ f'=0 \text{ when } 6x-6=0 \text{ this gives } x=1 \\ f''(x)=6>0 \text{ so this implies we have rel min at x=1 }\]

OpenStudy (anonymous):

omg that's what I have on my paper! so it is just like the other problem? nothing els?!

OpenStudy (freckles):

it was very similar to your previous problem

OpenStudy (anonymous):

cool! thanks!

OpenStudy (anonymous):

wait what do I do when I plug in the criticsl point and it equals zero??

OpenStudy (freckles):

when f''=0?

OpenStudy (anonymous):

oh no never mind lol I plugged it in wrong! but what are stationary points?

OpenStudy (freckles):

i think stationary point is another way to say critical point

OpenStudy (freckles):

also if x=a is a critical number and f''(a)=0, then this test is inclusive for relative max and min. It just tells us that the function is neither concave up or down at x=a.